# Multipoint distributions of the KPZ fixed point with compactly supported initial conditions

### v2

@@ 5.1 Density and Approximation
$+\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{+s\eta}\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(-\bm{\tau}^{\alpha,\beta}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}^{\alpha,\beta}_{+})\xi\right)\mathbf{1}_{\bm{\tau}^{\alpha,\beta}_{+}<\infty}\right]$
$-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(\bm{\tau}^{\alpha,\beta}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}^{\alpha,\beta}_{-})\eta\right)\mathbf{1}_{\bm{\tau}^{\alpha,\beta}_{-}>-\infty}\right]\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(-\bm{\tau}^{\alpha,\beta}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}^{\alpha,\beta}_{+})\xi\right)\mathbf{1}_{\bm{\tau}^{\alpha,\beta}_{+}<\infty}\right],$
@@
Applying similar arguments to the second and third term we see that
@@
$+\mathrm{e}^{\alpha\eta^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\widehat{\mathbf{B}}(\alpha)=s}\left[\exp\left(-\widehat{\bm{\tau}}_{+}\xi^{2}-\widehat{\mathbf{B}}(\widehat{\bm{\tau}}_{+})\xi\right)\mathbf{1}_{\widehat{\bm{\tau}}_{+}<\infty}\right$]
$-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\widehat{\mathbf{B}}(\alpha)=s}\left[\exp\left(\widehat{\bm{\tau}}_{-}\eta^{2}+\widehat{\mathbf{B}}(\widehat{\bm{\tau}}_{-})\eta-\widehat{\bm{\tau}}_{+}\xi^{2}-\widehat{\mathbf{B}}(\widehat{\bm{\tau}}_{+})\xi\right)\mathbf{1}_{|\widehat{\bm{\tau}}_{\pm}|<\infty}\right$],
@@
Hence by Proposition 5.2 we have
@@
$=\lim_{n\to\infty}\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}^{n}}(z_{1},\cdots,z_{m-1})$
$=\prod_{\ell=1}^{m-1}\oint\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1}).$
@@ 6 Reduction in the equal-time case
The goal of this section is to simplify the multipoint formula (1.6) under the additional assumption that all the time parameters $\tau_{1},\cdots,\tau_{m}$ are the same, say, equal to $1$. We will show in Section 6.1 that one can get rid of the additional contour integrals with respect to the parameters $z_{i}$’s in (1.6), and get a genuine Fredholm determinant formula for the equal-time multipoint distribution. To the best of our knowledge this Fredholm determinant formula is not written in the literature. Then in Section 6.2 we will show that our Fredholm determinant formula is equivalent to the path integral formula in [MQR21].
@@ 6.1 A new Fredholm determinant formula for the equal-time multipoint distribution of the KPZ fixed point
The following minor modification of [LO25, Proposition 2.1] holds for the KPZ fixed point with a general initial condition $\mathfrak{h}$ that is compactly supported.
@@
Let $\alpha_{1}<\cdots<\alpha_{m}$. Consider the KPZ fixed point starting from initial condition $\mathfrak{h}\in\mathrm{UC}_{c}$. Then the following formula for the multipoint distribution at space-time points $(\alpha_{1},1),\cdots,(\alpha_{m},1)$ holds:
$\mathbb{P}\left(\bigcap_{\ell=1}^{m}\mathcal{H}(\alpha_{\ell},1;\mathfrak{h})\leq\beta_{\ell}\right)=\det(\mathrm{I}+\mathbf{T}_{\mathfrak{h}})_{L^{2}(\{1,\cdots,m\}\times\Gamma_{1,\mathrm{R}})},$ (6.1)
where the operator $\mathbf{T}_{\mathfrak{h}}:L^{2}(\{1,\cdots,m\}\times\Gamma_{1,\mathrm{R}})\to L^{2}(\{1,\cdots,m\}\times\Gamma_{1,\mathrm{R}})$ has the following kernel:
@@
where $\mathbf{n}=(n_{1},\cdots,n_{m})\in\mathbb{Z}_{+}^{m}$ and
@@
$\cdot\prod_{i=1}^{m}\det\left[h_{i}(\xi_{a}^{(i)},\eta_{b}^{(i)})\right]_{a,b=1}^{k_{i}}\cdot\prod_{i=1}^{m}\prod_{\ell_{i}=1}^{k_{i}}\frac{1}{\mathrm{f}_{i}(\eta_{\ell_{i}}^{(i)})}.$
@@
Apply a generalized Andreief’s identity (see, e.g., [LO25, Lemma 1.2]) to the $\xi$-integrals, we get
$\widehat{\mathrm{D}}^{\mathbf{(n)}}_{\mathfrak{h}}$ $=\prod_{i=1}^{m}\frac{(n_{i}!)^{2}}{k_{i}!}\left(\prod_{i=1}^{m}\prod_{\ell_{i}=1}^{k_{i}}\int_{\Gamma_{1,\mathrm{R}}}\frac{\mathop{}\!\mathrm{d}\eta_{\ell_{i}}^{(i)}}{2\pi\mathrm{i}}\right)\det\left[\int_{1,\mathrm{L}}\frac{\mathop{}\!\mathrm{d}\xi_{1}}{2\pi\mathrm{i}}h_{i}(\xi_{1},\eta_{\ell_{i}}^{(i)})\cdot\frac{\chi_{\mathfrak{h}}(\eta_{\ell_{j}}^{(j)},\xi_{1})}{\mathrm{f}_{j}(\eta_{\ell_{j}}^{(j)})}\right]_{(i,\ell_{i}),(j,\ell_{j})}$ (6.7)
@@
where $\mathbf{T}_{\mathfrak{h}}(i,\zeta;j,\eta)$ is defined as in (6.2). Thus we have
$\mathbb{P}\left(\bigcap_{\ell=1}^{m}\mathcal{H}(\alpha_{\ell},1;\mathfrak{h})\leq\beta_{\ell}\right)$ $=\sum_{k_{1},\cdots,k_{m}\geq 0}\frac{1}{k_{1}!\cdots k_{m}!}\left(\prod_{i=1}^{m}\prod_{\ell_{i}=1}^{k_{i}}\int_{\Gamma_{1,\mathrm{R}}}\frac{\mathop{}\!\mathrm{d}\eta_{\ell_{i}}^{(i)}}{2\pi\mathrm{i}}\right)\det\left[\mathbf{T}_{\mathfrak{h}}(i,\eta_{\ell_{i}}^{(i)};j,\eta_{\ell_{j}}^{(j)})\right]_{(i,\ell_{i}),(j,\ell_{j})}$
$=\det(\mathrm{I}+\mathbf{T}_{\mathfrak{h}})_{L^{2}(\{1,\cdots,m\}\times\Gamma_{1,\mathrm{R}})}.$
@@ 6.2 Equivalence with the path integral formula of [MQR21]
The goal of this section is to prove that our new equal-time multipoint formula (6.1) is equivalent to the known ones. We will show that it is equivalent to the following path integral formula obtained in [MQR21]:
@@
Assume $\mathfrak{h}\in\mathrm{UC}_{c}$. Let $\mathbf{S}_{\mathfrak{h}}:=-\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}+\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\mathbf{1}_{\leq\beta_{2}}\cdots\mathbf{1}_{\leq\beta_{m}}\mathrm{e}^{(\alpha_{m}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}$. Then we have the following contour integral representation for the kernel of $\mathbf{S}_{\mathfrak{h}}$:
@@
$\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\cdots\mathbf{1}_{\leq\beta_{m}}\mathrm{e}^{(\alpha_{m}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}$ $=\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\cdots\mathbf{1}_{\leq\beta_{m-1}}\mathrm{e}^{(\alpha_{m-1}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}$
$-\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\cdots\mathbf{1}_{\leq\beta_{m-1}}\mathrm{e}^{(\alpha_{m-1}-\alpha_{m})}\mathbf{1}_{>\beta_{m}}\mathrm{e}^{(\alpha_{m}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})},$
@@
Repeating this argument for $\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\cdots\mathbf{1}_{\leq\beta_{i}}\mathrm{e}^{(\alpha_{i}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}$ with $i=m-1,\cdots,2$, we see that
@@
$S_{t,x}(p,q)=S_{t,x}^{*}(q,p)=\int_{\Gamma_{1,\mathrm{R}}}\frac{\mathop{}\!\mathrm{d}\eta}{2\pi\mathrm{i}}\mathrm{e}^{\frac{t}{3}\eta^{3}+x\eta^{2}+(p-q)\eta}=\int_{\Gamma_{1,\mathrm{L}}}\frac{\mathop{}\!\mathrm{d}\xi}{2\pi\mathrm{i}}\mathrm{e}^{-\frac{t}{3}\xi^{3}+x\xi^{2}-(p-q)\xi},$ (6.12)
@@
where the change of order of integration is justified by Proposition 2.3. A similar computation for the other two terms imply that
@@
where $\mathrm{F}_{i}(\zeta)=:\mathrm{e}^{(\alpha_{i}-\alpha_{i-1})\zeta^{2}+(\beta_{i}-\beta_{i-1})\zeta}$ and $c_{2}>c_{3}$. Similarly for any $2\leq i\leq m$ we have
@@
Combining (6.18) with (6.10), we arrive at the desired expression (6.9) for $\mathbf{S}_{\mathfrak{h}}(\lambda,\mu)$. ∎
@@
for any $\xi_{1}\in\Gamma_{1,\mathrm{L}}$ and $\xi_{2}\in\Gamma_{2,\mathrm{L}}^{\mathrm{in}}$. Now we write $\mathbf{T}_{\mathfrak{h}}:=L_{1}L_{2}$, where $L_{1}:L^{2}(\mathbb{R})\to L^{2}(\{1,\cdots,m\}\times\Gamma_{1,\mathrm{R}})$ has the following kernel:
@@
and $L_{2}:L^{2}(\{1,\cdots,m\}\times\Gamma_{1,\mathrm{R}})\to L^{2}(\mathbb{R})$ has the following kernel:
@@
$\det(\mathrm{I}+\mathbf{T}_{\mathfrak{h}})_{L^{2}(\{1,\cdots,m\}\times\Gamma_{1,\mathrm{R}})}=\det(\mathrm{I}+L_{1}L_{2})=\det(\mathrm{I}+L_{2}L_{1}):=\det(\mathrm{I}+\widehat{\mathbf{S}}_{\mathfrak{h}})_{L^{2}(\mathbb{R})},$
@@
$+\sum_{i=2}^{m}\int_{\Gamma_{1,\mathrm{R}}}\frac{\mathop{}\!\mathrm{d}\eta}{2\pi\mathrm{i}}\left(\prod_{\ell=1}^{i}\int_{\Gamma_{\ell,\mathrm{L}}^{\mathrm{in}}}\frac{\mathop{}\!\mathrm{d}\xi_{\ell}}{2\pi\mathrm{i}}\right)\frac{\prod_{\ell=1}^{i}\mathrm{F}_{\ell}(\xi_{\ell})}{\prod_{\ell=2}^{i-1}(\xi_{\ell}-\xi_{\ell+1})\cdot(\xi_{i}-\eta)}\frac{\mathrm{e}^{\lambda\xi_{1}-\mu\xi_{2}}}{\mathrm{f}_{i}(\eta)}\cdot\chi_{\mathfrak{h}}(\eta,\xi_{1})\cdot\mathbf{1}_{\mu\leq 0}.$
Comparing (6.21) with (6.9) we see that $\widehat{\mathbf{S}}_{\mathfrak{h}}(\lambda,\mu)=\mathbf{S}_{\mathfrak{h}}(\mu+\beta_{1},\lambda+\beta_{1})$. Thus
$\det(\mathrm{I}+\mathbf{S}_{\mathfrak{h}})_{L^{2}(\mathbb{R})}=\det(\mathrm{I}+\widehat{\mathbf{S}}_{\mathfrak{h}})_{L^{2}(\mathbb{R})}=\det(\mathrm{I}+\mathbf{T}_{\mathfrak{h}})_{L^{2}(\{1,\cdots,m\}\times\Gamma_{1,\mathrm{R}})}.$

### v3

@@ 5.1 Density and Approximation
$\quad+\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathrm{e}^{+s\eta}\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(-\bm{\tau}^{\alpha,\beta}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}^{\alpha,\beta}_{+})\xi\right)\mathbf{1}_{\bm{\tau}^{\alpha,\beta}_{+}<\infty}\right]$
$\quad-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(\bm{\tau}^{\alpha,\beta}_{-}\eta^{2}+\mathbf{B}(\bm{\tau}^{\alpha,\beta}_{-})\eta\right)\mathbf{1}_{\bm{\tau}^{\alpha,\beta}_{-}>-\infty}\right]$
$\quad\qquad\quad\cdot\mathbb{E}_{\mathbf{B}(0)=s}\left[\exp\left(-\bm{\tau}^{\alpha,\beta}_{+}\xi^{2}-\mathbf{B}(\bm{\tau}^{\alpha,\beta}_{+})\xi\right)\mathbf{1}_{\bm{\tau}^{\alpha,\beta}_{+}<\infty}\right],$
@@
By applying similar arguments to the second and third term, we see that
@@
$\quad+\mathrm{e}^{\alpha\eta^{2}}\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\widehat{\mathbf{B}}(\alpha)=s}\left[\exp\left(-\widehat{\bm{\tau}}_{+}\xi^{2}-\widehat{\mathbf{B}}(\widehat{\bm{\tau}}_{+})\xi\right)\mathbf{1}_{\widehat{\bm{\tau}}_{+}<\infty}\right$]
$\quad-\int_{\mathbb{R}}\mathop{}\!\mathrm{d}s\,\mathbb{E}_{\widehat{\mathbf{B}}(\alpha)=s}\left[\exp\left(\widehat{\bm{\tau}}_{-}\eta^{2}+\widehat{\mathbf{B}}(\widehat{\bm{\tau}}_{-})\eta-\widehat{\bm{\tau}}_{+}\xi^{2}-\widehat{\mathbf{B}}(\widehat{\bm{\tau}}_{+})\xi\right)\mathbf{1}_{|\widehat{\bm{\tau}}_{\pm}|<\infty}\right$],
@@
Hence, by Proposition 5.2 we have
@@
$=\lim_{n\to\infty}\prod_{\ell=1}^{m-1}\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}^{n}}(z_{1},\ldots,z_{m-1})$
$=\prod_{\ell=1}^{m-1}\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{\ell}}{2\pi\mathrm{i}z_{\ell}(1-z_{\ell})}\mathrm{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1}).$
@@ 6 Reduction in the equal-time case
The goal of this section is to simplify the multipoint formula (1.6) under the additional assumption that all the time parameters $\tau_{1},\ldots,\tau_{m}$ are the same, say, equal to $1$. We will show in Section 6.1 that one can get rid of the additional contour integrals with respect to the parameters $z_{i}$’s in (1.6), and get a Fredholm determinant formula for the equal-time multipoint distribution. Then in Section 6.2, we will show that the Fredholm determinant formula is equivalent to the path integral formula in [MQR21].
@@ 6.1 A new Fredholm determinant formula for the equal-time multipoint distribution of the KPZ fixed point
The following Fredholm determinant formula holds for the KPZ fixed point with a general initial condition $\mathfrak{h}$ that is compactly supported. We remark that an analogous version for the narrow wedge initial condition was obtained in [LO25, Proposition 2.1], with a slight modification on the operator space.
@@
Let $\alpha_{1}<\cdots<\alpha_{m}$. Consider the KPZ fixed point starting from initial condition $\mathfrak{h}\in\mathrm{UC}_{c}$. Then the following formula for the multipoint distribution at space-time points $(\alpha_{1},1),\ldots,(\alpha_{m},1)$ holds:
$\mathbb{P}\left(\bigcap_{\ell=1}^{m}\mathcal{H}(\alpha_{\ell},1;\mathfrak{h})\leq\beta_{\ell}\right)=\det(\mathrm{I}+\mathbf{T}_{\mathfrak{h}})_{L^{2}(\{1,\ldots,m\}\times\Gamma_{1,\mathrm{R}})},$ (6.1)
where the operator $\mathbf{T}_{\mathfrak{h}}:L^{2}(\{1,\ldots,m\}\times\Gamma_{1,\mathrm{R}})\to L^{2}(\{1,\ldots,m\}\times\Gamma_{1,\mathrm{R}})$ has the following kernel:
@@
where $\mathbf{n}=(n_{1},\ldots,n_{m})\in\mathbb{Z}_{+}^{m}$ and
@@
$\quad\cdot\prod_{i=1}^{m}\det\left[h_{i}(\xi_{a}^{(i)},\eta_{b}^{(i)})\right]_{a,b=1}^{k_{i}}\cdot\prod_{i=1}^{m}\prod_{\ell_{i}=1}^{k_{i}}\frac{1}{\mathrm{f}_{i}(\eta_{\ell_{i}}^{(i)})}.$
@@
We apply a generalization of the Andreief’s identity obtained in [LO25, Lemma 1.2] to the $\xi$-integrals. It gives
$\widehat{\mathrm{D}}^{\mathbf{(n)}}_{\mathfrak{h}}$ $=\prod_{i=1}^{m}\frac{(n_{i}!)^{2}}{k_{i}!}\left(\prod_{i=1}^{m}\prod_{\ell_{i}=1}^{k_{i}}\int_{\Gamma_{1,\mathrm{R}}}\frac{\mathop{}\!\mathrm{d}\eta_{\ell_{i}}^{(i)}}{2\pi\mathrm{i}}\right)\det\left[\int_{\Gamma_{1,\mathrm{L}}}\frac{\mathop{}\!\mathrm{d}\xi_{1}}{2\pi\mathrm{i}}h_{i}(\xi_{1},\eta_{\ell_{i}}^{(i)})\cdot\frac{\chi_{\mathfrak{h}}(\eta_{\ell_{j}}^{(j)},\xi_{1})}{\mathrm{f}_{j}(\eta_{\ell_{j}}^{(j)})}\right]_{(i,\ell_{i}),(j,\ell_{j})}$ (6.7)
@@
where $\mathbf{T}_{\mathfrak{h}}(i,\zeta;j,\eta)$ is defined as in (6.2). Thus, we have
$\mathbb{P}\left(\bigcap_{\ell=1}^{m}\mathcal{H}(\alpha_{\ell},1;\mathfrak{h})\leq\beta_{\ell}\right)$ $=\sum_{k_{1},\ldots,k_{m}\geq 0}\frac{1}{k_{1}!\cdots k_{m}!}\left(\prod_{i=1}^{m}\prod_{\ell_{i}=1}^{k_{i}}\int_{\Gamma_{1,\mathrm{R}}}\frac{\mathop{}\!\mathrm{d}\eta_{\ell_{i}}^{(i)}}{2\pi\mathrm{i}}\right)\det\left[\mathbf{T}_{\mathfrak{h}}(i,\eta_{\ell_{i}}^{(i)};j,\eta_{\ell_{j}}^{(j)})\right]_{(i,\ell_{i}),(j,\ell_{j})}$
$=\det(\mathrm{I}+\mathbf{T}_{\mathfrak{h}})_{L^{2}(\{1,\ldots,m\}\times\Gamma_{1,\mathrm{R}})}.$
@@ 6.2 Equivalence with the path integral formula of [MQR21]
The goal of this section is to prove that the equal-time multipoint formula (6.1) is equivalent to the following path integral formula obtained in [MQR21] when the initial condition is compactly supported.
@@
Assume $\mathfrak{h}\in\mathrm{UC}_{c}$. Let
$\mathbf{S}_{\mathfrak{h}}:=-\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}+\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\mathbf{1}_{\leq\beta_{2}}\cdots\mathbf{1}_{\leq\beta_{m}}\mathrm{e}^{(\alpha_{m}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}.$
Then we have the following contour integral representation for the kernel of $\mathbf{S}_{\mathfrak{h}}$:
@@
$\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\cdots\mathbf{1}_{\leq\beta_{m}}\mathrm{e}^{(\alpha_{m}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}$
$=\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\cdots\mathbf{1}_{\leq\beta_{m-1}}\mathrm{e}^{(\alpha_{m-1}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}$
$\quad-\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\cdots\mathbf{1}_{\leq\beta_{m-1}}\mathrm{e}^{(\alpha_{m-1}-\alpha_{m})}\mathbf{1}_{>\beta_{m}}\mathrm{e}^{(\alpha_{m}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})},$
@@
Repeating this argument for $\mathbf{1}_{\leq\beta_{1}}\mathrm{e}^{(\alpha_{1}-\alpha_{2})\partial^{2}}\cdots\mathbf{1}_{\leq\beta_{i}}\mathrm{e}^{(\alpha_{i}-\alpha_{1})\partial^{2}}\mathbf{K}_{1,\alpha_{1}}^{\mathrm{hypo}(\mathfrak{h})}$ with $i=m-1,\ldots,2$, we see that
@@
$S_{t,x}(p,q)=S_{t,x}^{*}(q,p):=\int_{\Gamma_{1,\mathrm{R}}}\frac{\mathop{}\!\mathrm{d}\eta}{2\pi\mathrm{i}}\mathrm{e}^{\frac{t}{3}\eta^{3}+x\eta^{2}+(p-q)\eta}=\int_{\Gamma_{1,\mathrm{L}}}\frac{\mathop{}\!\mathrm{d}\xi}{2\pi\mathrm{i}}\mathrm{e}^{-\frac{t}{3}\xi^{3}+x\xi^{2}-(p-q)\xi},$ (6.12)
@@
where the change of order of integration is justified by Proposition 2.3. A similar computation for the other two terms in (6.11) implies that
@@
where $\mathrm{F}_{i}(\zeta)=:\mathrm{e}^{(\alpha_{i}-\alpha_{i-1})\zeta^{2}+(\beta_{i}-\beta_{i-1})\zeta}$ and $c_{2}>c_{3}$. Similarly, for any $2\leq i\leq m$ we have
@@
By combining (6.18) with (6.10), we arrive at the desired expression (6.9) for $\mathbf{S}_{\mathfrak{h}}(\lambda,\mu)$. ∎
@@
for any $\xi_{1}\in\Gamma_{1,\mathrm{L}}$ and $\xi_{2}\in\Gamma_{2,\mathrm{L}}^{\mathrm{in}}$. Now we write $\mathbf{T}_{\mathfrak{h}}:=L_{1}L_{2}$, where $L_{1}:L^{2}(\mathbb{R})\to L^{2}(\{1,\ldots,m\}\times\Gamma_{1,\mathrm{R}})$ has the following kernel:
@@
and $L_{2}:L^{2}(\{1,\ldots,m\}\times\Gamma_{1,\mathrm{R}})\to L^{2}(\mathbb{R})$ has the following kernel:
@@
$\det(\mathrm{I}+\mathbf{T}_{\mathfrak{h}})_{L^{2}(\{1,\ldots,m\}\times\Gamma_{1,\mathrm{R}})}=\det(\mathrm{I}+L_{1}L_{2})=\det(\mathrm{I}+L_{2}L_{1}):=\det(\mathrm{I}+\widehat{\mathbf{S}}_{\mathfrak{h}})_{L^{2}(\mathbb{R})},$
@@
$\quad+\sum_{i=2}^{m}\int_{\Gamma_{1,\mathrm{R}}}\frac{\mathop{}\!\mathrm{d}\eta}{2\pi\mathrm{i}}\left(\prod_{\ell=1}^{i}\int_{\Gamma_{\ell,\mathrm{L}}^{\mathrm{in}}}\frac{\mathop{}\!\mathrm{d}\xi_{\ell}}{2\pi\mathrm{i}}\right)\frac{\prod_{\ell=1}^{i}\mathrm{F}_{\ell}(\xi_{\ell})}{\prod_{\ell=2}^{i-1}(\xi_{\ell}-\xi_{\ell+1})\cdot(\xi_{i}-\eta)}\frac{\mathrm{e}^{\lambda\xi_{1}-\mu\xi_{2}}}{\mathrm{f}_{i}(\eta)}\cdot\chi_{\mathfrak{h}}(\eta,\xi_{1})\cdot\mathbf{1}_{\mu\leq 0}.$
Comparing (6.21) with (6.9), we see that $\widehat{\mathbf{S}}_{\mathfrak{h}}(\lambda,\mu)=\mathbf{S}_{\mathfrak{h}}(\mu+\beta_{1},\lambda+\beta_{1})$. Thus
$\det(\mathrm{I}+\mathbf{S}_{\mathfrak{h}})_{L^{2}(\mathbb{R})}=\det(\mathrm{I}+\widehat{\mathbf{S}}_{\mathfrak{h}})_{L^{2}(\mathbb{R})}=\det(\mathrm{I}+\mathbf{T}_{\mathfrak{h}})_{L^{2}(\{1,\ldots,m\}\times\Gamma_{1,\mathrm{R}})}.$